# There Was Never a Bottleneck in Concept Bottleneck Models

### v1

@@ 3.3 Concept Bottleneck Models
Figure 1: Graphical models of the different systems described. We consider two concepts and two-dimensional representations. We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $p_{\theta}(\bm{z}|\bm{x})$ (black arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $q_{\phi}(\bm{y}|\bm{x})$ (green arrows). Concept Bottleneck Models try to make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs (ours) try to make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $q(z_{j}|c_{j})$ (red arrows).
4 On the Lack of a Bottleneck in CBMs
As explained in Section 1, one of the main advantages commonly attributed to CBMs over vanilla models is their ability to perform concept-level interventions and propagate the effects to the model’s predictions. For instance, consider a scenario in which we aim to estimate $p(\hat{\bm{y}}\mid c_{j}=\alpha,\bm{x})$—that is, the model’s predicted output $\hat{\bm{y}}$ had the concept $c_{j}$ been set to the value $\alpha$. In CBMs, this intervention is typically performed through the latent representation $z_{j}$, according to:
[…]
However, in the original formulation of CBMs (described in Section 3.3), the conditional distribution $p(z_{j}\mid c_{j})$ is not defined—there is no directed path from $c_{j}$ to $z_{j}$ in Figure 1(c). Intuitively, because $z_{j}$ may encode information about $\bm{x}$ beyond $c_{j}$, it cannot be fully determined by $c_{j}$ alone. This leads to a key question: how can interventions be performed in CBMs if $p(z_{j}\mid c_{j})$ is unknown? In what follows, we address this question and argue that current practices often violate basic Bayesian principles, leading to inconsistencies and limiting the architectural flexibility of the model. To resolve this, we introduce Minimal CBMs (MCBMs), a theoretically grounded framework that addresses these issues and enables principled interventions.
4.1 Intervention Mechanisms in Concept Bottleneck Models
As previously discussed, the conditional distribution $p(z_{j}\mid c_{j})$ is not explicitly defined in standard CBMs. To enable a mechanism for performing interventions, two constraints are typically imposed:
  1. (i)
    Concepts $c_{j}$, for $j=1,2,\dots,m$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [38]).
  2. (ii)
Given that $\sigma$ is an invertible function, this setup permits defining $p(z_{j}\mid c_{j})=\sigma^{-1}(c_{j})$. However, since $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$, this inversion becomes ill-defined at the binary extremes. To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are approximated as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9]. Nevertheless, this practice introduces two issues:
  1. 1.
    One-vs-Rest strategies exhibit the following limitations: (i) individual binary classifiers tend to be biased toward the negative class, (ii) multiple classifiers may simultaneously assign a "positive" label to the same input, and (iii) the predicted probabilities across classifiers are typically uncalibrated [39]. These limitations are further illustrated in Section 5.3.
  2. 2.
    The intervention procedure based on the sigmoid inverse does not adhere to Bayes’ rule, i.e., $p(z_{j}\mid c_{j})=\frac{q_{\phi}(c_{j}\mid z_{j})p(z_{j})}{p(c_{j})}$. While $p(c_{j})$ can be approximated using its empirical distribution, the marginal $p(z_{j})$ remains unknown and is challenging to estimate.
4.2 Minimal Concept Bottleneck Models
As previously discussed, CBMs lack an explicit mechanism for enforcing a bottleneck, which often undermines their primary advantages. To address this limitation, we introduce Minimal Concept Bottleneck Models (MCBMs), which explicitly impose an Information Bottleneck. This ensures that each representation $z_{j}$ not only retains all information about the associated concept $c_{j}$, but also contains information only about $c_{j}$. In other words, $z_{j}$ becomes a minimal sufficient representation of $c_{j}$. To achieve this, we introduce a representation head $q_{\phi}(\hat{z}j\mid c_{j})$ that predicts $z_{j}$ from $c_{j}$. This encourages $z_{j}$ to discard information unrelated to $c_{j}$, as doing so improves the predictive accuracy of $q_{\phi}(\hat{z}_{j}\mid c_{j})$. Formally, this objective corresponds to minimizing the conditional mutual information $I(Z_{j};X\mid C_{j})$. We note that if $z_{j}$ is a representation of $\bm{x}$, then the following propositions are equivalent: (i) $I(Z_{j};X|C_{j})=0$, (ii) the Markov Chain $X\leftrightarrow C_{j}\leftrightarrow Z_{j}$ is satisfied and (iii) $p(z_{j}|c_{j})=p(z_{j}|x)$. To minimize $I(Z_{j};X\mid C_{j})$, we leverage the following identity, which is proven in Appendix B.3:
Figure 1(d) illustrates how MCBMs extend CBMs by introducing the representation head $q(\hat{z}_{j}\mid c_{j})$, depicted with red arrows. These models are trained to jointly optimize the objectives given in Equations (1), (2), and (4). Although the KL Divergence in Equation (4) does not admit a closed-form solution in general, we show in Appendix B.4 that, under the distributional assumptions listed in Table 1, it reduces to the mean squared error between $f_{\theta}(x)$ and $g^{z}_{\phi}(c_{j})$.
4.3 Practical Considerations for Optimizing MCBMs
To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations. First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution described in Section 3.1, replacing integrals with summations over the dataset. Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [40], allowing for low-variance gradient estimates during backpropagation. $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big{|}f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 1), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$. Combining the considerations above yields the final objective for training MCBMs, as shown in Equation (5), where $\beta$ and $\gamma$ are hyperparameters. The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs. s. Detailed training algorithms for the various cases listed in Table 1 are provided in Appendix C.
4.4 Intervention Mechanisms in Minimal Concept Bottleneck Models
As discussed at the beginning of this section, intervention procedures require access to $p(z_{j}\mid c_{j})$. As outlined in Section 4.1, this dependence anchors the intervention process to specific architectural design choices in CBMs. Moreover, such procedures violate Bayesian principles, as the prior distribution $p(z_{j})$ is typically unknown. In MCBMs, since $z_{j}$ is explicitly designed to contain information only about $c_{j}$, modifying $z_{j}$ corresponds to intervening solely on $c_{j}$. This is reflected in the graphical model illustrated in Figure 1(d), where MCBMs introduce a directed path from $c_{j}$ to $z_{j}$ through the intermediate variable $\hat{z}_{j}$. This structure permits the computation of:
[…]
which simplifies to $p(z_{j}\mid c_{j})=q_{\phi}(z_{j}\mid c_{j})$ when $p(z_{j}\mid\hat{z}_{j})=\delta(z_{j}-\hat{z}_{j})$, which holds when the objective in Equation (4) is optimized to convergence—or, equivalently, when $z_{j}$ contains information exclusively about $c_{j}$. As shown in Table 1, we define the encoder distribution as $q_{\phi}(z_{j}\mid c_{j})=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I)$. The mean function $g^{z}_{\phi}(c_{j})$ is chosen according to the following rules:
  1. (i)
  2. (ii)
    For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$. This mirrors the Prototypical Learning [41] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed, as adopted in [42].
  3. (iii)
    For continuous concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot c_{j}$.
Here, $\lambda$ is a scaling constant that controls the norm of the latent representation and is fixed to $\lambda=3$ across all experiments. Regarding the variance term, we set: (i) $\sigma_{x}=0$in the case of CBMs to obtain a deterministic encoder in line with their original formulation, and (ii) $\sigma_{x}=\sigma_{\hat{z}}=1$for MCBMs.
@@ 5 Experiments
In this section, we present a series of experiments designed to empirically demonstrate that CBMs fail to impose an effective bottleneck, even in simple settings. We also examine the consequences of this limitation. In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence. Additional implementation details, including encoder architectures and training hyperparameters for each experiment, are provided in Appendix D.
@@ 5.1 Do CBMs and MCBMs Leak Information?
In some works, it is assumed that the task $\bm{y}$ is fully defined by the concepts $\bm{c}$. However, it is unrealistic to have a finite set of concepts that we understand as humans and that totally describe an arbitrarily complex task. In general, some of the nuisances $\bm{n}$ that describe the input $\bm{x}$ will also influence $\bm{y}$. Specifically, we decompose the nuisances as $\bm{n}=\{\bm{n}_{y},\bm{n}_{\bar{y}}\}$, where $\bm{n}_{y}$ captures nuisances that, together with the concepts, describe $\bm{y}$, while $\bm{n}_{\bar{y}}$ comprises those that are independent of $\bm{y}$ (see Figure 2(a)). As shown through this text, CBMs have no incentive to remove information that is not related to the concepts. Additionally, since the objective of models in general is to solve $\bm{y}$, they have incentives to preserve not only $\bm{c}$ but also $\bm{n}_{y}$ in their representations. Although this may lead to improved performance in solving the task, it comes at the expense of interpretability and the ability to intervene in the system. For example if $z_{j}\in\bm{z}$ were intended to represent the concept $c_{j}\in\bm{c}$, then one could believe that an intervention is made solely on $c_{j}$ by modifying $z_{j}$. However $z_{j}$ could also contain information on $\bm{n}_{y}$ (or even $\bm{n}_{\bar{y}}$, as we see later), so we would be intervening also the nuisances by modifying $z_{j}$, thereby invalidating any causal conclusion drawn. By introducing the Information Bottleneck in equation (4), we force the model to remove the nuisances from $z_{j}$, thus restoring the validity of any causal analysis about $c_{j}$. We note that this will necessarily negatively impact the performance in solving the task $\bm{y}$ by the nature of the problem. If $\bm{c}$ is incomplete and information from $\bm{n}$ is necessary to solve $\bm{y}$ (i.e., $\bm{n}_{\bar{y}}\neq\emptyset$), the performance will decrease if $\bm{z}$ does not contain information about $\bm{n}$. We analyze next i $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representation for different versions of CBMs and datasets.
@@
CIFAR-10 MPI3D Shapes3D CelebA CUB
Vanilla $33.5\pm 4.3$ $35.0\pm 1.9$ $45.5\pm 6.4$ $10.7\pm 3.5$ $3.8\pm 1.0$
CBM $34.7\pm 0.2$ $28.1\pm 0.5$ $8.1\pm 2.9$ $11.5\pm 3.5$ $3.8\pm 0.8$
CEM $35.1\pm 0.1$ $43.2\pm 5.2$ $15.8\pm 3.9$ $16.8\pm 3.3$ $3.9\pm 1.1$
MCBM (low $\gamma$) $24.8\pm 0.6$ $10.7\pm 0.1$ $2.5\pm 0.1$ $4.3\pm 0.0$ $3.4\pm 0.9$
MCBM (medium $\gamma$) $19.1\pm 1.0$ $6.7\pm 0.2$ $0.2\pm 0.3$ $\bm{7.9\pm 2.9}$ $2.8\pm 0.8$
MCBM (high $\gamma$) $\bm{11.0\pm 0.5}$ $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $8.0\pm 2.4$ $\bm{2.4\pm 1.0}$
@@
(a) CBM ($\beta=1$)
(b) MCBM ($\beta=1$, $\gamma=5$)
Figure 3: Average prediction $\hat{y}$ in animal (0) vs. vehicle (1) classification for the different classes when excluding airplane and bird from the concept set.
(c) CBM ($\beta=1$)
(d) MCBM ($\beta=1$, $\gamma=5$)
Figure 4: Average prediction $\hat{y}$ for animal (purple) vs. vehicle (yellow) when intervening on $\bm{z}$ to match the x-axis class instead of the true class (y-axis).
We analyze whether $\bm{n}_{y}$ is encoded in $\bm{z}$—that is, whether the dependency illustrated by the orange arrow in Figure 2(b) is present—and examine its consequences. To illustrate this, we design a simple experiment based on the CIFAR-10 dataset [43]. Since all classes in CIFAR-10 correspond to either animals or vehicles, we define the target $\bm{y}$ as the binary task of distinguishing between these two categories. The concepts $\bm{c}$ are defined using all classes except airplane and bird. Under a proper bottleneck, the model should be unable to accurately predict $\bm{y}$ for these excluded classes, as their information is not captured by any concept. Figure 4 shows the mean predictions $\hat{y}$ for each class, comparing a CBM and an MCBM. The CBM confidently predicts $\bm{y}$ even for the excluded classes, indicating a substantial degree of information leakage. In contrast, the MCBM significantly mitigates this issue, displaying low confidence for excluded classes while maintaining high confidence for those included in the concept set. To assess the impact on interventions, Figure 4 shows model predictions after simulating an intervention corresponding to each concept. In the CBM, predictions following interventions exhibit reduced confidence for certain classes—an effect attributed to residual nuisance information in the representation. This effect is notably reduced in the MCBM.
To systematically analyze this phenomenon, we consider several datasets and define distinct task–concept configurations for each. Specifically: (i) In CIFAR-10 [43], we adopt the described scenario: $\bm{y}$ corresponds to a binary classification between animals and vehicles, $\bm{n}_{y}$ includes airplane and bird, and $\bm{c}$ represents the remaining classes. (ii) In MPI3D [44], $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}_{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors. (iii) In Shapes3D [30], $\bm{y}$ is the shape, $\bm{n}_{y}$ includes the floor color and wall color, $\bm{n}_{\bar{y}}$ is the orientation, and $\bm{c}$ includes the remaining factors. (iv) In CelebA [45], $\bm{y}$ is gender, $\bm{c}$ consists of eight randomly selected attributes, and $\bm{n}_{y}$ includes the remaining factors, all of which are correlated with $\bm{y}$. (v) In CUB [46], $\bm{y}$ corresponds to the bird species, $\bm{c}$ includes the concepts from seven randomly selected attribute groups, and $\bm{n}_{y}$ comprises the attributes from the remaining 20 groups. Further details for each setup are provided in Appendix D. We recall that, while our MCBMs natively support multiclass concepts, the other models evaluated in this work are limited to binary concepts. Therefore, to ensure fair comparisons, factors with $m$ classes are represented as $m$ separate binary concepts.
To assess the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we aim to estimate $I(N_{j};Z\mid C)$—that is, the amount of additional information that $\bm{z}$ provides about $n_{j}$ beyond what is already captured by $\bm{c}$. Since this quantity is intractable in practice, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$. Here, $h_{\psi}^{c}$ and $h_{\psi}^{cz}$ are MLP classifiers trained to predict $n_{j}$ from $\bm{c}$ and $(\bm{c},\bm{z})$, respectively. In Table 2, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we refer to as the Uncertainty Reduction Ratio (URR). From these results, we draw the following conclusions: (i) regular CBMs do not consistently remove more nuisance information than VMs—in fact, contrary to common belief, CBMs can sometimes impose a weaker bottleneck than VMs; (ii) CEMs representations are those with more nuisances, i.e., they always introduce a weaker bottleneck even than VMs; and (iii) MCBMs are the models removing more nuisances from the representation, especially when $\gamma$ increases.
@@
Table 3: Average URR for task-unrelated nuisances.
@@
MCBM (low $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (medium $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
On the one hand, neural networks typically discard information from the input, as their layers are generally non-invertible [47, 48]. On the other hand, there is no incentive to retain $\bm{n}_{\bar{y}}$ in the representation $\bm{z}$, since it is irrelevant for predicting $\bm{y}$. From this, one might expect that $\bm{z}$ should not contain information about $\bm{n}_{\bar{y}}$. However, prior work has shown that neural representations often retain information that is not directly relevant to the target task [24, 49]. To investigate this phenomenon, we analyze if $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings in our experimental setup where $\bm{n}_{\bar{y}}$ is non-empty. We report the URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in Table 3, and observe the following: (i) in VMs, CBMs, and CEMs, nuisance variables remain present in the representation, even when they are irrelevant to the task $\bm{y}$; and (ii) MCBMs consistently eliminate task-irrelevant nuisances across all values of $\gamma$, which is expected: since there is no incentive to preserve $\bm{n}_{\hat{y}}$, and MCBMs explicitly encourage its removal, such information tends to be discarded.

### v2

@@ 2.4 Minimal Concept Bottleneck Models
Figure 2: Graphical models of the different systems described for two concepts and two-dimensional representations. We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows). Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
3 Related Work
Information Leakage in Concept Bottleneck Models
Information leakage occurs when the learned representation $\bm{z}$ encodes information outside the concept set $\bm{c}$, reducing both interpretability and intervenability [10, 11]. This arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios [18]. Concept Embedding Models (CEMs) [19] were proposed to mitigate the accuracy–interpretability trade-off in CBMs. However, this trade-off is fundamentally limited by the chosen concept set, and CEMs may even be less interpretable than standard CBMs, as their more entropic representations tend to amplify information leakage. Recently, this critique has been further formalized: Parisini et al. [20] proposed information-theoretic benchmarks for assessing interpretability and Makonnen et al. [21] introduced new metrics grounded in usable mutual information [22]. Havasi et al. [18] also introduced Hard Concept Bottleneck Models (HCBMs), which predict $\bm{y}$ from binarized concept predictions $\hat{c}_{j}$ rather than from $\bm{z}$ (see Appendix D), thereby imposing an ad-hoc Information Bottleneck. Extensions such as Autoregressive CBMs and Stochastic CBMs [18, 23] incorporate dependencies between concepts.
Information Bottleneck in Representation Learning
The Information Bottleneck (IB) [12] provides a principled way to balance preserving information about a factor with compressing the representation. In this framework, a representation is sufficient if it retains all the information about the variable of interest, and minimal if it contains only that information [24, 25, 26]. Computation of these quantities is intractable, so variational methods have been developed to derive tractable evidence bounds [27, 28].
@@ 4 Experiments
In this section, we present a series of experiments designed to empirically demonstrate that CBMs fail to impose an effective bottleneck, even in simple settings. We also examine the consequences of this limitation. In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence. Further implementation details—including encoder architectures and training hyperparameters for each experiment—are provided in Appendix E and at https://github.com/antonioalmudevar/minimal_cbm.
@@ 4.1 Do CBMs and MCBMs Leak Information?
Some works assume that the task $\bm{y}$ is fully defined by the concepts $\bm{c}$. However, it is unrealistic to expect a finite set of human-understandable concepts to completely describe an arbitrarily complex task. In practice, certain nuisances $\bm{n}$ that describe the input $\bm{x}$ also influence $\bm{y}$. Specifically, we decompose the nuisances as $\bm{n}=\{\bm{n}_{y},\bm{n}_{\bar{y}}\}$, where $\bm{n}_{y}$ captures nuisances that, together with the concepts, describe $\bm{y}$, while $\bm{n}_{\bar{y}}$ comprises those independent of $\bm{y}$ (see Figure 3(a)). As discussed throughout this work, CBMs provide no incentive to remove information unrelated to the concepts. Moreover, since the objective of most models is to solve $\bm{y}$, they are incentivized to preserve not only $\bm{c}$ but also $\bm{n}_{y}$. While this may improve task performance, it comes at the expense of interpretability and valid interventions. For instance, if $z_{j}\in\bm{z}$ is intended to represent the concept $c_{j}\in\bm{c}$, , one might assume that modifying $z_{j}$ corresponds solely to intervening on $c_{j}$. However, if $z_{j}$ also encodes information about $\bm{n}_{y}$ (or even $\bm{n}_{\bar{y}}$ as we show later), then modifying $z_{j}$ also affects the nuisances, invalidating any causal conclusions. By introducing the Information Bottleneck in Equation 4, we explicitly constrain the model to remove nuisances from $z_{j}$, thereby restoring the validity of causal analyses involving $c_{j}$. We note, however, that this necessarily reduces task performance: if $\bm{c}$ is incomplete and solving $\bm{y}$ requires information from $\bm{n}$ (i.e., $\bm{n}_{y}\neq\emptyset$), excluding $\bm{n}$ from $\bm{z}$ will lower predictive performance.
We next examine whether $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representations across different CBM variants and datasets. For this purpose, we define the following task–concept configurations: (i) MPI3D[29], where $\bm{y}$ is the object shape, $\bm{n}_{y}$ the horizontal axis, $\bm{n}_{\bar{y}}$ the vertical axis, and $\bm{c}$ the remaining generative factors; (ii) Shapes3D[30], where $\bm{y}$ is the shape, $\bm{n}_{y}$ the floor color and wall color, $\bm{n}_{\bar{y}}$ the orientation, and $\bm{c}$ the remaining factors; (iii) CIFAR-10[31], where $\bm{y}$ is the standard classification task, $\bm{c}$ consists of 64 of the 143 attributes extracted by Oikarinen et al. [32] using GPT-3 [33], and $\bm{n}_{y}$ the remaining attributes, since all nuisances are correlated with $\bm{y}$; and (iv) CUB[34], where $\bm{y}$ is the bird species, $\bm{c}$ includes concepts from twelve randomly selected attribute groups, and $\bm{n}_{y}$ the attributes from the remaining 20 groups. While MCBMs natively support multiclass concepts, the other baselines in our study are limited to binary concepts. To ensure fairness, factors with $k$ classes are therefore represented as $k$ binary concepts.
@@
MPI3D Shapes3D CIFAR-10 CUB
Vanilla $35.0\pm 1.9$ $45.5\pm 6.4$ $19.8\pm 0.7$ $3.8\pm 1.0$
CBM [9] $28.1\pm 0.5$ $18.1\pm 2.9$ $18.5\pm 0.7$ $3.8\pm 0.8$
CEM [19] $43.2\pm 5.2$ $15.8\pm 3.9$ $27.2\pm 0.8$ $3.9\pm 1.1$
ARCBM [18] $28.2\pm 1.7$ $18.4\pm 2.1$ $18.2\pm 0.6$ $3.9\pm 0.9$
SCBM [23] $24.3\pm 0.4$ $21.8\pm 1.4$ $18.3\pm 0.7$ $3.6\pm 0.8$
MCBM (low $\gamma$) $10.7\pm 0.1$ $2.5\pm 0.1$ $18.0\pm 0.5$ $3.4\pm 0.9$
MCBM (medium $\gamma$) $6.7\pm 0.2$ $0.2\pm 0.3$ $18.1\pm 0.5$ $2.8\pm 0.8$
MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $\bm{17.6\pm 0.5}$ $\bm{2.4\pm 1.0}$
@@
To measure the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we estimate $I(N_{j};Z\mid C)$—the information about $n_{j}$ contained in $\bm{z}$ beyond what is explained by $\bm{c}$. Since this quantity is intractable, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$. Here, $h_{\psi}^{c}$ and $h_{\psi}^{cz}$ are MLP classifiers trained to predict $n_{j}$ from $\bm{c}$ and $(\bm{c},\bm{z})$, respectively. In Table 3, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we call Uncertainty Reduction Ratio (URR). From these results, we conclude that: (i) CBMs tend to reduce nuisance information compared to VMs, though not consistently; (ii) CEMs generally preserve the largest amount of nuisance information, often exceeding even VMs; (iii) ARCBMs and HCBMs show no systematic advantage over CBMs in terms of nuisance removal; and (iv) MCBMs provide the strongest reduction of nuisance information, particularly as $\gamma$ increases, which enforces a stricter bottleneck.
@@
Table 4: Average value of URR for task-unrelated nuisances.
@@
ARCBM $8.7\pm 1.4$ $26.6\pm 0.5$
SCBM $7.0\pm 0.3$ $21.7\pm 1.7$
MCBM (l $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (m $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
MCBM (h $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$
Neural networks typically discard input information, as their layers are non-invertible [35, 36]. Moreover, since $\bm{n}_{\bar{y}}$ is irrelevant for predicting $\bm{y}$, there is no incentive to retain it in the representation $\bm{z}$. One might therefore expect $\bm{z}$ to be free of $\bm{n}_{\bar{y}}$. However, prior work has shown that neural representations often preserve information not directly related to the task [24, 37]. To examine this, we analyze whether $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings where $\bm{n}_{\bar{y}}$ is non-empty. Table 4 reports URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in. We find that: (i) as in Table 3, CEMs retain more nuisance information than even VMs; and (ii) MCBMs are the only models that consistently eliminate nuisances across all values of $\gamma$, as expected: with no incentive to preserve $\bm{n}_{\bar{y}}$ and an explicit penalty for doing so, such information is naturally discarded.